OpenStudy (anonymous):

find the area bounded by the given curves , the y-axis and the given values of y.

OpenStudy (anonymous):

y = (x+2)(x-1)(x-3) when y=0, x= -2 , 1 , 3 So, the area is bounded by lines x=-2 , x=3 , and the curve and x-axis . y= x^3 - 2x^2 - 5x +6 Area bounded = ∫ y dx Integrate from x=-2 to x=3 =∫x^3 - 2x^2 - 5x +6 = [ x^4/4 - 2x^3/3 - 5x^2/2 + 6x ] x=-1 to x=3 = 125/12

OpenStudy (anonymous):

x= sqrt (a^2 - y^2)

OpenStudy (anonymous):

from y=0 to y=a ....a is constant